# RClicks: Realistic Click Simulation for Benchmarking Interactive Segmentation

### v1

@@ Appendix E Datasheet for the Benchmark
Datasheets for datasets datasheet facilitate communication between data creators and users in the form of the questionnaire explicating motivation, data acquisition process, and potential use cases. In this document, we provide a datasheet for the benchmark.
@@ E.1 Motivation
To address this, researchers often use a simple strategy to simulate user inputs. This strategy involves generating a single click for each interaction round by selecting the largest error region from the previous round and clicking at the furthest point from the boundaries of this region (center point). However, previous works [ritm, moskalenko2024tetris] have shown that assuming users click in the center of an object is overly simplistic and unrealistic.

### v2

@@ Appendix E Datasheet for the Benchmark
Datasheets for datasets Gebru et al. [2021] facilitate communication between data creators and users in the form of the questionnaire explicating motivation, data acquisition process, and potential use cases. In this document, we provide a datasheet for the benchmark.
@@ E.1 Motivation
To address this, researchers often use a simple strategy to simulate user inputs. This strategy involves generating a single click for each interaction round by selecting the largest error region from the previous round and clicking at the furthest point from the boundaries of this region (center point). However, previous works [Sofiiuk et al., 2022, Moskalenko et al., 2024] have shown that assuming users click in the center of an object is overly simplistic and unrealistic.